OpenStudy (beccaboo333):

If you were to use the substitution method to solve the following system, choose the new equation after the expression equivalent to y from the first equation is substituted into the second equation. 8x - y = -9 4x - 3y = -22

OpenStudy (beccaboo333):

A. 4(-8x - 9) - 3y = -22 B. 4x - 3(-8x - 9) = -22 C. 4(8x + 9) - 3y = -22 D. 4x - 3(8x + 9) = -22

OpenStudy (anonymous):

D

OpenStudy (beccaboo333):

are you sure Rod?

OpenStudy (anonymous):

I EXPLAIN IT

OpenStudy (beccaboo333):

ALRIGHT

OpenStudy (anonymous):

8x - y = -9 -8x -8x -y=-8x-9 -1 <-y=-8x-9 (-8x-9) -1 -1 -1 y=8x+9

OpenStudy (anonymous):

4x-3(8x+9)=-22 so did u understand it becca

OpenStudy (beccaboo333):

yes slighty. thank you. help me with a few more? :3

OpenStudy (anonymous):

ok becca i try

OpenStudy (beccaboo333):

thanks Rod :D